# Sotopia-RL: Reward Design for Social Intelligence

### v1

@@ C.5 Data Constent
Appendix D AI Assistant Details
@@
Appendix E Additional Experimental Results
E.1 Full Main Results
Table 5 shows the comprehensive evaluation results for our method, including the evaluation of all the evaluation dimensions in Sotopia-eval on Sotopia-hard and Sotopia-all.
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Overall
Sotopia-hard
BC 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
BC + SR 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.81 1.84 4.14 0.00 -0.09 0.31 5.61 2.95
Singular Goal 9.00 2.74 4.93 -0.04 -0.05 0.61 6.64 3.41
Scaled Goal 8.94 1.82 3.83 -0.04 -0.01 0.76 6.74 3.15
Direct Rel 8.96 3.61 4.92 -0.01 -0.11 0.59 7.24 3.60
Direct Kno 8.99 2.56 6.06 0.00 -0.01 0.75 6.93 3.61
Direct Goal 8.99 2.49 4.94 -0.00 -0.06 0.91 7.21 3.49
Direct Goal +Kno +Rel 9.01 3.41 5.53 -0.26 -0.06 1.16 7.81 3.80
Sotopia-all
BC 8.99 3.08 4.56 -0.09 -0.06 0.57 7.80 3.55
BC + SR 8.98 2.52 4.19 -0.06 -0.06 0.57 7.36 3.36
Uniform Goal 8.87 2.49 4.19 0.00 -0.02 0.44 6.76 3.25
Singular Goal 8.99 3.38 5.46 -0.07 -0.08 0.66 7.72 3.72
Scaled Goal 8.97 2.76 4.70 -0.12 -0.06 0.55 7.97 3.54
Direct Rel 8.98 3.95 5.54 -0.03 -0.05 0.65 8.33 3.91
Direct Kno 8.98 3.00 6.42 -0.03 -0.03 0.63 7.76 3.82
Direct Goal 8.99 3.11 5.74 -0.06 -0.06 0.76 8.11 3.80
Direct Goal +Kno +Rel 8.99 3.81 6.00 -0.61 -0.08 0.93 8.57 3.94
E.2 Full Ablation Results
Table 6: Ablation results on Sotopia-hard with different partner and evaluator models. Sotopia-RL represents the direct attribution + reward combination. Goal-RL represents the direct attribution + Goal-only reward. BC represents behavior cloning (Ziegler et al., 2020). Top block: partner model ablation (evaluator model fixed to GPT-4o). Bottom block: evaluator model ablation (partner model fixed to BC). Full results for Figure 7  7.
Partner Model Sotopia-RL Goal-RL BC
Goal Overall Goal Overall Goal Overall
BC 7.75 3.79 7.21 3.49 6.76 3.16
GPT-4o 7.39 3.69 6.57 3.35 5.91 3.04
Claude-3.7-Sonnet 6.71 3.24 6.61 3.43 6.54 3.35
Deepseek-v3 6.38 3.13 5.57 2.95 6.00 2.97
Qwen2.5-72B-Instruct 7.64 3.74 6.80 3.45 6.88 3.20
Evaluator Model Sotopia-RL Goal-RL BC
Goal Overall Goal Overall Goal Overall
GPT-4o 7.81 3.80 7.21 3.49 6.76 3.16
GPT-4 8.25 4.33 7.35 3.78 6.76 3.32
Claude-3.7-Sonnet 7.35 3.44 6.49 3.23 6.06 2.96
Deepseek-v3 7.75 4.02 7.05 3.65 6.83 3.35
Qwen2.5-72B-Instruct 8.05 4.16 7.49 3.61 7.19 3.26
E.3 Best-of-$N$ Evaluation Results
Attribution Dimension Bel Rel Kno Sec Soc Fin Goal Overall
BC 9.01 2.49 3.37 0.00 -0.06 0.56 6.76 3.16
BC + SR 8.99 2.41 3.66 0.00 -0.10 0.61 6.84 3.20
Uniform Goal 8.96 2.51 3.90 -0.06 -0.09 0.60 6.79 3.23
Singular Goal 8.99 2.56 3.72 -0.07 -0.09 0.62 7.07 3.26
Scaled Goal 8.99 2.58 4.02 -0.06 -0.06 0.77 7.20 3.35
Direct Rel 9.00 2.86 4.14 -0.05 0.00 0.45 6.83 3.32
Direct Kno 8.99 2.61 4.49 0.00 -0.12 0.58 7.04 3.37
Direct Goal 8.99 2.44 4.13 -0.05 -0.19 0.56 6.97 3.27
Direct Goal +Kno +Rel 9.00 2.48 3.86 -0.01 -0.13 0.68 6.85 3.25
E.4 Safety Evaluation Results
Model Benign Adversarial Benign Adversarial Benign Jailbreak
BC 0.61 0.77 0.78 0.92 0.86 0.39
Sotopia-RL 0.58 0.75 0.50 0.90 0.86 0.40
E.5 Diversity Evaluation Results
[…]
Table 9: Diversity evaluation results for different models. We calculate the turn number and word numbers for social interactions.
Model Avg. Turn Number Avg. Word Number
BC 14.44 37.17
Sotopia-$\pi$ 10.81 35.50
GOAL-RL 19.41 51.83
Sotopia-RL 19.59 76.53
Appendix F Additional Analysis of Utterance-level Reward
Correlation annotator1 annotator2 annotator3 annotator4 GPT-4o
annotator1 1.000 0.812 0.931 0.891 0.771
annotator2 0.812 1.000 0.737 0.717 0.636
annotator3 0.931 0.737 1.000 0.818 0.756
annotator4 0.891 0.717 0.818 1.000 0.664
GPT-4o 0.771 0.636 0.756 0.664 1.000
Figure 14: Reward attribution and reward combination examples. On the left, it explains 4 types of attribution methods (uniform, scaled, singular, and direct). On the right, it explains 4 types of different reward combination methods (Rel-RL, Goal-RL, Kno-RL, Sotopia-RL) where all of them are based on direct reward attribution and Sotopia-RL is the combined one. More details are in Appendix §G and §H.
Appendix G Additional Details about Reward Attribution
@@
G.1 Attribution Template
We prompt the attribution LLM using the template defined in ATTRIBUTION_TEMPLATE  G.1. Specifically, we populate the fields goal, agent_background, and conversation using the episode logs from Sotopia. Detailed descriptions and prompt for attribution_instruction and dimension_description are provided in  G.2  G.4 and Section H, respectively.
@@ G.1 Attribution Template
G.2 Direct Attribution
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use DIRECT_ATTRIBUTION prompt in  G.2.
@@ G.4 Singular Attribution
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use SINGULAR_ATTRIBUTION prompt in  G.4.
SINGULAR ATTRIBUTION 1. Input Context: You will receive the dialogue history between two conversational agents. You will also be provided with the social goal of one of the agents. 2. Objective: Identify the most critical utterance that has the highest impact on the final ga oal achievement, whether it is bad or good impact. Note, you should only consider how critical an utterance is to the final goal achievement, not the quality of the utterance itself. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. 3. Additional Guidelines: The conversation history will be given in a unique key of "Utterance utterance number by agent name" for each utterance. Please only return the key of the most critical utterance. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. Note: You will also be given a formatting instruction for instructions. Please follow the instruction to ensure the evaluation process runs smoothly.
G.5 Uniform Attribution
@@
Appendix H Additional Details about Reward Dimensions
@@
We supply dimension_description using the definitions of goalcompletion, relationship, and knowledge, adapted from prompts in Sotopia-eval. For goalcompletion, we include an additional explanation due to the ambiguity of the original definition. The right side of Figure 14 shows a concrete example for reward combination.
Goal Dimension Description Goal refers to the reiteration of the agent’s social goals and the analysis of their achievement. A higher score indicates significant progress or achievement of the stated goals, while a lower score indicates minimal or no progress.
DOMAIN SPECIFIC SCORING GUIDELINES: ! Note: The following scoring guidelines are specific to the domain of goal and should be used in conjunction with the domain-specific scale. The domain specific rules ultimately override the general scoring scale. Here are the specific rules:
- The highest score should be assigned to the utterance that is most relevant to the goal. In general, avoid assigning the highest score to more than one utterance unless they are equally critical. - A lower score should be assigned to the utterances that are not relevant to the goal or do not contribute to its achievement. The lowest score should be assigned to the utterances that do not make any progress towards the goal judging by the goal description and the conversation history. - A lower score should be assigned to the utterances that are not effective in achieving the goal, judging by the response of the other agent. Effective utterances are those that lead to a positive response from the other agent, while ineffective utterances are those that lead to a negative or neutral response. - Note that you should only consider the contribution to the goal achievement. For each utterance, assess whether the goal is achieved. If a goal is already achieved, the utterance should not be assigned a score higher than 1.
@@
Appendix I Additional Case Study
Figure 15,16,17,18 show four examples of conversation between the model trained with behavior cloning (BC) and the model trained with Sotopia-RL (Sotopia-RL). After RL-based training, Sotopia-RL is able to find collaborative solutions in many scenarios and has multi-turn strategies to achieve its social goals in specific cases.

### v3

@@ F.5 Data Constent
Appendix G The Use of Large Language Models (LLMs)
@@
Appendix H Additional Details about Reward Attribution
@@
H.1 Attribution Template
We prompt the attribution LLM using the template defined in ATTRIBUTION_TEMPLATE  H.1. Specifically, we populate the fields goal, agent_background, and conversation using the episode logs from Sotopia. Detailed descriptions and prompt for attribution_instruction and dimension_description are provided in  H.2  H.4 and Section I, respectively.
@@ H.1 Attribution Template
H.2 Direct Attribution
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use DIRECT_ATTRIBUTION prompt in  H.2.
@@ H.4 Singular Attribution
To generate the attribution_instruction field within the ATTRIBUTION_TEMPLATE, we use SINGULAR_ATTRIBUTION prompt in  H.4.
SINGULAR ATTRIBUTION 1. Input Context: You will receive the dialogue history between two conversational agents. You will also be provided with the social goal of one of the agents. 2. Objective: Identify the most critical utterance that has the highest impact on the final ga oal achievement, whether it is bad or good impact. Note, you should only consider how critical an utterance is to the final goal achievement, not the quality of the utterance itself. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. 3. Additional Guidelines: The conversation history will be given in a unique key of ”Utterance utterance number by agent name” for each utterance. Please only return the key of the most critical utterance. Consider both the individual utterance and the responses from the other agent, as both affect the final outcome. Note: You will also be given a formatting instruction for instructions. Please follow the instruction to ensure the evaluation process runs smoothly.
H.5 Uniform Attribution
@@
Appendix I Additional Details about Reward Dimensions
We conducted a more fine-grained experiment of different weights of reward dimensions.
@@ Appendix I Additional Details about Reward Dimensions
We supply dimension_description using the definitions of goalcompletion, relationship, and knowledge, adapted from prompts in Sotopia. For goalcompletion, we include an additional explanation due to the ambiguity of the original definition. The right side of Figure 10 shows a concrete example for reward aggregation.
Goal Dimension Description Goal refers to the reiteration of the agent’s social goals and the analysis of their achievement. A higher score indicates significant progress or achievement of the stated goals, while a lower score indicates minimal or no progress. DOMAIN SPECIFIC SCORING GUIDELINES: ! Note: The following scoring guidelines are specific to the domain of goal and should be used in conjunction with the domain-specific scale. The domain specific rules ultimately override the general scoring scale. Here are the specific rules: - The highest score should be assigned to the utterance that is most relevant to the goal. In general, avoid assigning the highest score to more than one utterance unless they are equally critical. - A lower score should be assigned to the utterances that are not relevant to the goal or do not contribute to its achievement. The lowest score should be assigned to the utterances that do not make any progress towards the goal judging by the goal description and the conversation history. - A lower score should be assigned to the utterances that are not effective in achieving the goal, judging by the response of the other agent. Effective utterances are those that lead to a positive response from the other agent, while ineffective utterances are those that lead to a negative or neutral response. - Note that you should only consider the contribution to the goal achievement. For each utterance, assess whether the goal is achieved. If a goal is already achieved, the utterance should not be assigned a score higher than 1.
@@
Appendix J Case Study
Based on Figure 14, we analyze a multi-turn interaction where Naomi (Sotopia-RL model) persuades Sophia (behavior cloning model) to share the only blanket on a cold camping night. This example illustrates how our social agent trained under multi-objective RL is able to generate a single utterance that advances multiple social dimensions at once. For instance, “I know it feels good to …” both pursues Naomi’s goal and strengthens the relationship bond (Rel). Furthermore, “I can always …, and I’ll stay close to the flames to ensure I stay cozy” conveys her willingness to adapt using external knowledge (Kno) while remaining considerate. Finally, “Let’s try sharing the blanket.” explicitly states her goal, aligning with the goal completion dimension (Goal). Together, these utterances show how the trained agent integrates goal pursuit with friendliness and informativeness. Additional case studies are provided in Appendix §J.
Additionally, Figure 15,16,17,18 show four examples of conversation between the model trained with behavior cloning (BC) and the model trained with Sotopia-RL (Sotopia-RL). After RL-based training, Sotopia-RL is able to find collaborative solutions in many scenarios and has multi-turn strategies to achieve its social goals in specific cases.